OpenStudy (anonymous):

What is the equation of parabola that opens right, has a vertex at origin and a vertical stretch of 1/2?

OpenStudy (anonymous):

what does "vertical stretch of 1/2" mean?

OpenStudy (anonymous):

nvm can u help me with this one. Find the sum of x and y-coordinates of the vertex of the parabola given by equation 4x + 20 = (y - 7)2

OpenStudy (anonymous):

(y-7) to the power of 2

OpenStudy (anonymous):

if you write it in this form: 4a(x-h) = (y-k)^2 (h, k) is your vertex.

OpenStudy (anonymous):

so factor out that 4 on the left: 4(x + 5) = (y - 7)^2 now tell me what is (h, k) ?

OpenStudy (anonymous):

(-5,7)

OpenStudy (anonymous):

good. that is your vertex... now you can answer the queston... :)

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

great... :)

OpenStudy (anonymous):

one more question please

OpenStudy (anonymous):

What is the equation of the axis of symmetry of the parabola y2 - 3x - 10y + 13 = 0?